TestBeamOct2018




This section will summarize the main documents, informations and results of the October 2018 testbeam. A fiber based telescope was installed in the testbeam area (Prévessins site).

  • The layout is presented below.
  • The data can be found on the LPHE cluster under (TO BE CONFIRMED).

Open issues

Installation

  • Compressed air: at the installation (SPS), we used nitrogen connection from VELO (Raphael) and RICH (Didier). This is however not enough (flux). We stay with this configuration.

  • Inspection patrol: grounding is not good (the item profiles have insulating layer). Was fixed on 13.10.

DAQ Logbook

  • 10.10: first DAQ basic with noise: very high noise is seen in a few channels, even without bias. Due to temperature of the STiCs? Grounding?
  • 15.10: Guido changed the fpga firmware so that data are not accepted beyond a delay of 00FF after the trigger word. New firmware (v74) loaded in : cdfpga/FPGA_firmware/current_tb2/set_fpga_files
  • 16.10: new firmware version from Guido (v77) filters the data too early and too late.
  • 16.10: agreed with the other groups to have lower intensity for 1h. Alex (phone: 162259) was contacted and explained how to change the beam configuration. Procedure in the logbook.

Run plan

This chapter focuses on the measurement planned and the status.

1. Experiment [Waiting for discussion]

Telescope mechanics and read-out

Detector layout

detector_layout.png

The telescope is based on scintillating fibers coupled to SiPMs. There are 4 SiPMs per board, each of them is coupled to 2 STiCs (2x64 channels). In total, there are 8 STiCs per board, connected to the FPGA.

SIPMs

This section should have a table of each layer. This should include detail such as breakdown voltage - dead channel (SiPM), name, physical number (STiC), ip number(STiC).

Detector inventory

detector_inventory.png

SiPMs features details

sipm_detail.png

Installed boards

Initial configuration : 10.Oct.2018

Board IP Position Dead channels
25 MASTER 5 -
17 8 -
27 11 -
08 2 -
05 20 stic7 ch26
21 14 stic3 ch11-13
12 17 stic6 ch1-11-12
24 23 stic3 ch12-13
Replacement boards

Stored at CERN

recalibration and configuration are done on those boards

Board IP # of STiCs attached issues length
16 4 STiC 5 ch0 dead SHORT
10 4 STiC 5 ch5,24 bad SHORT
06 4 STiC 6 ch0,7 dead SHORT
03 2 no issues LONG

Stored at EPFL

Board IP # of STiCs attached issues length
13 2

STiC 4 ch6,8,13 dead

STiC 5 ch5,7,13,15,31,33 dead

LONG
04 2

STiC 7 issues with TTH calibration

STiC 5 ch25 bad

SHORT
29 4 modified by Guido SHORT
NB: issues mentioned in tables above concern the STiCs attached to the boards during recalibration

Trigger

Trigger to Hit delay map

Trigger to Hit delay for the different layers (Master and Slave levels)

Level Board position Delay LPHE Delay SPS Delay PS
Master 5 192 ns 243 ns 155 ns
Level 1 Slaves 8, 11, 20 239 ns 291 ns 194 ns
Level 2 Slaves 2, 14, 17 279 ns 331 ns 233 ns
Level 2 Slave (long cable) 23 290 ns 343 ns 245 ns

Data acquisition

Machines and network

CAEN crate

IP address setup. This section should have IP number for setting up at CERN

DAQ machine

Data acquisition

daq_setup.png

ETH scan

  • (if needed : spi_stic/stic_support.c, in function do_eth_scan_verif_v2, line 837 : change freq_eth_adjust to 130Hz)
  • cdfpga/config_tb2 : remplacer le DAQ par DAQ_all_stic_en to enable all the STiCs.
  • ./ source runETH_verif to launch a scan on all STiCs
  • follow the scan with screen : screen -ls and screen -r job#
  • run iteratively the scan to converge to right value
Remark : if ETh scan is bad, always put back a good backup before leaving, so that the next operator can work without worries.

RUNS (DAQ)

DATE RUN NUMBER NUMBER OF TRIGGERS RATE OV (V) ENABLED STICS COMMENTS
13/10 1019 185k High 5.5 CONFIG1* BOARD11 had 123k triggers
13/10 1021 281k High 4.5 CONFIG1* ok
13/10 1022 295k High 3.5 CONFIG1* ok
14/10 2010 270k High 3.5 ? ok
15/10 1027 3M High 3.5 CONFIG1*  
15/10 1028 1M High 3.5 CONFIG1* trig cnt a bit different in the log file
15/10 1044 1.2M <=60kHz 4.5 CONFIG1* profit from re-start of beam after end of access
15/10 1045 4.2M 60kHz 4.5 CONFIG1*  
15/10 1046 1.2M 60kHz 5.5 CONFIG1*  
15/10 1047 4M 60kHz 5.5 CONFIG1* Board positions 8, 14, 17 : underflow frown
15/10 1048 4M 60kHz 5.5 CONFIG1* Board position 14 : underflow frown
15/10 1049 4M 60kHz 5.5 CONFIG1* Board position 17 : underflow frown
16/10 4008 100k <=30kHz 3.5 CONFIG2* No data from board IP12 - firmware version 74
16/10 4009 120k 30kHz 3.5 CONFIG2* No data from board IP12 - firmware version 74
16/10 4010 383k 30kHz <= f <= 100kHz 3.5 CONFIG2* No data from board IP12 - firmware version 74
* CONFIG1
stics_enabled_config1.png

RUNS PS

DATE Coordinator RUN # Local Run # RPC TOF Calo Beam intensity # TRIGGERS RATE OV Firmware STiCS Comments
18/10 2 4015 y y y -5 GeV focused on the calo 4K (at 1:14 scale) ~10 kHz (5000 ev/0.6 s) 3.5 v77 all STiCs enabled (changed by hand)  
18/10 ? 5001 y y y '' 20K '' 3.5 v77 'all STiCs enabled  
18/10 4 5002 y y y '' 100K '' 3.5 v77 all STiCs enabled  
18/10 4 5003 y y y '' 38K '' 3.5 v77 all STiCs enabled  
18/10 5 5004 y y y '' 174k '' 4.5 v77 all STiCs enabled  
20/10   7000 n y y -5 GeV 100K '' 5.5 v77 all STiCs enabled  
20/10   7001 n y y -5 GeV 5K '' 4.5 v77 all STiCs enabled

quadrupoles chaged at:

QDE6:-201

QDO7:-199

20/10   7002 n y y -8GeV   '' 3.5 v77 all STiCs enabled

quadrupoles chaged at:

QDE6:-334

QDO7:-318

CRATE turned off at some point of the run. Data to be checked.

21/10   8000 y y y     '' 4.5 v77 all STiCs enabled

Mastergate not working anymore.

--> have to suppress first trigger afterwards

21/10   8001 (dummy) y y n

-8GeV

(ZT9A.ZT9-EXP.186)

  '' 4.5

v77

all STiCs enabled

''

triggers realigned with the beamin x and y

QDE6: -322

QFO7: -296

BHZ3:-191

21/10   8002 y y n

+10 Gev

positive paricles (e+, mu+, p)

  '' 4.5 v77 all STiCs enabled

Qdri focused at 9.5m

402

371

412

388

21/10  

8003

(dummy)

y y n

+10GeV

positive paricles (e+, mu+, p)

 

1000 TRIG/ spill (0.6 s)

= 1.67 kHz

4.5 v77 all STiCs enabled

playing with PLL of STIC 2,3 of board 27

Qdri focused at 7.0 m

QDE6: 412.3

QFO7: 388.8

21/10   8004 y y n

+10 GeV

positive paricles (e+, mu+, p)

25k events

run continued up to 100K

1.67 kHz

5 spill/cycle

4.5 v77 all STiCs enabled

Qdri focused at 7.0 m

QDE6: 412.3

QFO7: 388.8

21/10   8005 y y n

+10 GeV

positive paricles (e+, mu+, p)

access@75K

100k

idem 3.5 v77 all STiCs enabled idem
21/10   8006 y y n

+10 GeV

positive paricles (e+, mu+, p)

access@29K

100K

idem 5.5 v77 all STiCs enabled idem
21/10   8007 y y n

+10GeV

positive paricles (e+, mu+, p)

100K idem 4.5 v77 all STiCs enabled idem
    8008 y y n

+10GeV

positive paricles (e+, mu+, p)

100K idem 3.5 v77 all STiCs enabled idem
PS: Shower characterisation for SHiP calorimetry
  • Trigger: Cherenkov counter, set on electrons.
  • Beam: -5 GeV electrons, focused on the SciFi. Configuration file: ZT9A.ZT9-EXP.233
  • Rate: ~50 electrons / spill
  • Dead time in the trigger chain disabled.
  • Distance between closest iron plate and first SciFi layer: 320mm. Distance between iron plates: 25mm.
  • Radiation length of iron: ~1.7mm. So, 1 iron plate (4mm) represents ~0.25*X0.

DATE Local Run # # iron plates Thickness # TRIGGERS OV Firmware STiCs Comments
19/10 6001 0 0 2k 4.5 v77 all Pressure in the Cherenkov rose, it seems we were not only triggering on electrons. Beam >5cm wide, Walter shrinks the beam.
19/10 6002 0 0 2.5k 4.5 v77 all Pressure in the Cherenkov set. Beam cannot be shrunk. We'll have to use the small trigger in front.
19/10 6005 4 1*X0 5k 4.5 v77 all  
19/10 6006 8 2*X0 5.3k 4.5 v77 all  
19/10 6007 16 4*X0 5.7k 4.5 v77 all  
19/10 6008 24 6*X0 2.8k 4.5 v77 all  
19/10 6010 24 6*X0 5k 4.5 v77 all  
Data Visualization

Analysis

[AFTER TAKING DATA, it is needed to be validated - HOW ?]

RawData bakup (done on 21.10.2018)

In Federico's public SHIP lxplus (username: fredish)

fredish@lxplus7NOSPAMPLEASE.cern.ch:/afs/cern.ch/work/f/fredish/public/october-testbeam-scifi/

-- SurapatEkIn1 - 2018-10-08-- SurapatEkIn1 - 2018-10-08-- SurapatEkIn1 - 2018-10-08-- SurapatEkIn1 - 2018-10-08

plane # layer # board pos # CAEN ch # board # board IP IP stic card # stic chip # old name group name detector # Operational Voltage (@3.5) Operational Voltage (@4.5) Output Fast Control IP
Topic attachments
I Attachment History Action Size Date Who Comment
PDFpdf LY_Matv3.pdf r1 manage 61.5 K 2018-10-12 - 17:05 SurapatEkIn1 Light yield measurement on SciFi mats
PDFpdf LY_SF36_23_Absorber3mm_posAM_Summary.pdf r1 manage 199.2 K 2018-10-12 - 17:05 SurapatEkIn1 Light yield measurement on SciFi mats
PNGpng LightYieldPerMat_TCor1.png r1 manage 70.0 K 2018-10-12 - 17:05 SurapatEkIn1 Light yield measurement on SciFi mats
PNGpng Screen_Shot_2018-10-13_at_23.40.05.png r1 manage 131.4 K 2018-10-13 - 23:41 AnaCavalcante config1_stics_enabled
PNGpng TWIKI.png r1 manage 80.2 K 2018-10-14 - 11:32 SoniaAminaBouchiba  
PNGpng daq_setup.png r1 manage 81.7 K 2018-10-14 - 13:08 SoniaAminaBouchiba  
PNGpng detector_inventory.png r1 manage 80.0 K 2018-10-14 - 11:35 SoniaAminaBouchiba SiPMs, Stic number, overvoltage etc
PNGpng detector_layout.png r1 manage 63.5 K 2018-10-12 - 17:10 AdeleLuceCaroleMackowiak position of the layers and boards in the detector
PNGpng sipm_detail.png r1 manage 77.4 K 2018-10-14 - 12:02 SoniaAminaBouchiba  
PNGpng stics_enabled_config1.png r1 manage 131.4 K 2018-10-13 - 23:42 AnaCavalcante  
Edit | Attach | Watch | Print version | History: r46 < r45 < r44 < r43 < r42 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r46 - 2018-11-27 - BlakeLeverington
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback